Two airplanes leave Dallas at the same time and fly in opposite directions. One airplane travels 80 miles per hour faster then the other. After 3 hours they are 2940 miles apart. What is rhe rate of each airplane?
Answer by checkley75(3666)

3*X+3(X+80)=2940
3X+3X+240=2940
6X=2940-240
6X=2700
X=2700/6
X=450 MPH FOR THE SLOW PLANE
450+80=530 MPH FOR THE FASTER PLANE.
PROOF
3*450+3*530=2940
1350+1590=2940
2940=2940